OpenStudy (anonymous):

A museum had 380 people sign up for e-mails. if 4/5 of these people were from Miami, how many were NOT from Miami

OpenStudy (anonymous):

1/5 or 380* 1/5

OpenStudy (anonymous):

how to calculate the answer: need to divide 380 through five, can't see but can put into parts: 380 = 350 + 30 and 350/5 = 70, 30/5 = 6 so 380/5 = 76 people were NOT from Miami

OpenStudy (anonymous):

thank you